# Fixing Code Generation Errors for Large Language Models

### v1

@@
Abstract
@@
Code generation leverages artificial intelligence technologies, particularly Large Language Models (LLMs), to automatically produce source code, enhancing software development efficiency and reducing repetitive tasks. However, the LLMs’ generated code often fails to pass test cases and requires substantial human effort to fix errors. Previous studies focused on better prompts or improving LLMs’ capability but ignored why LLMs failed. In this paper, we first reproduced 14 LLMs, including GPT-3.5-turbo and 13 open-source LLMs, on the HumanEval dataset. We extracted 12,837 code generation errors and conducted an in-depth analysis of their causes, which led to the identification of 19 distinct error causes. Our empirical analysis indicated that three of these causes can be directly fixed. Consequently, we proposed a fixing method called LlmFix, which addresses these three types of errors through a three-step process: filtering code for indentation correction, truncating redundant generated code, and importing missing modules. Experimental results demonstrate that LlmFix can fix these three types of errors, significantly improving the performance of 14 LLMs on HumanEval and MBPP datasets with average increases of 9.5% and 5.4%, respectively.
Index Terms:
@@
Large Language Model, Code Generation, Automatic Fixing
I Introduction
Code generation refers to the process of converting natural language descriptions of software development tasks into automatically generated code snippets that effectively solve these specific tasks [1, 2]. In real-world applications, developers expect the generated code to not only compile successfully but also perform the intended tasks accurately, in alignment with specified requirements or problem statements. Effective code generation can significantly reduce developers’ time and effort in coding, enabling them to concentrate on more complex and innovative aspects of software engineering [3, 4].
Large Language Models (LLMs) based on Transformer architecture have become powerful tools for various natural language processing (NLP) tasks, including sentiment analysis [5, 6] and language translation [7]. LLMs pre-trained on extensive codebases and fine-tuned specifically for code generation have also demonstrated promising performance in generating code [2, 8]. Furthermore, Zheng et al. [9] conducted a comprehensive survey of existing open-source code generation LLMs, reporting their performance on the widely used HumanEval dataset.
Despite advancements in LLMs for code generation, their output often contains errors, prompting researchers to explore various methods for improvement. Liu et al. [10] evaluated ChatGPT’s capabilities for two code generation tasks, including text-to-code and code-to-code generation. They designed prompts by leveraging the chain-of-thought strategy with multi-step optimizations. Their findings demonstrated that carefully crafted prompts for ChatGPT can significantly enhance the performance of code generation. White et al. [11] explored how to optimize interactions with LLMs by designing efficient prompt patterns, highlighting the importance of crafting effective prompts. They also provided concrete implementation examples and demonstrated the practical effects of these patterns. Chen et al. [12] formalized an algorithm for learning from natural language feedback at training time, which they call imitation learning from language feedback.
While previous studies improve code generation performance by refining prompts or enhancing the model itself, they do not investigate the root causes of test case failures in model-generated code, nor do they explore possible methods to fix these errors directly after the generation. However, understanding and fixing these errors is crucial to achieving efficient code generation.
In this paper, we first conduct an empirical study to investigate the performance of LLMs in code generation, and identify the types of errors they generate. Specifically, we select 14 LLMs and, referencing the experimental results of Zheng et al. [9] on the HumanEval dataset [13], accurately reproduce their code generation performance on the HumanEval dataset, followed by a series of in-depth empirical studies. Our findings reveal that LLM performance generally correlates positively with parameter scale, while inference time shows no significant association with parameter size. Then, to investigate the reasons behind LLMs failing in code generation tasks, we add “try-except” code in each generated code and collect 12,837 errors related to 14 types of exceptions (e.g., AssertionError, NameError, etc.). According to in-depth manual analysis, we ultimately identify 19 specific causes leading to these errors. Our research reveals that 16 causes (e.g., Inappropriate Argument, Unbalanced Delimiters, etc.) belong to the incapability of LLMs, and only three causes were amenable to direct automated fixing including: 1) Inconsistent Indentation, the code fails to pass the compilation due to wrong indentation; 2) Function Overflow, the completed function is appended many redundant and partial code as LLM does not stop in time; 3) Missing Import, the generated code invokes modules without import.
@@ I Introduction
Based on these empirical research findings, we design LlmFix, a method specifically tailored to automatically fix errors caused by the above three causes, which includes the following three corresponding operations: 1) Code Filtering, searching for inconsistent indentation and replacing it with uniform tab-based indentation. 2) Code Truncation, discarding redundant code located after the function completion. 3) Importing Missing Modules, automatically adding corresponding import statements at the beginning of the code when detecting the use of a module that has not been imported. We evaluate our LlmFix on two datasets (i.e., HumanEval and MBPP). Experimental results demonstrate that LlmFix can fix these three types of errors which are amenable for automated fixing and improve the performance of the 14 LLMs by an average of 9.5% and 5.4% on two code evaluation benchmarks, respectively.
@@
    Comprehensively investigating the code generation performance of 14 LLMs, analyzing the relationship between model performance, inference time, and parameter size.
@@
    Systematically categorizing and analyzing 12,837 code generation errors to reveal error distributions and reasons.
@@
    Proposing LlmFix for automatically fixing code generation errors and demonstrating its effectiveness of 14 LLMs on HumanEval and MBPP datasets.
@@
    Sharing our replication package [14].
This paper is structured as follows. Section II introduces the related work on code generation. Section III provides the empirical study on LLM replication and error analysis. Section IV presents our error fixing method LlmFix, which is evaluated in Section V. Section VI provides implications for developers and researchers and the threats to the validity of our study. Section VII summarizes this paper.
II Related Work
This section describes the existing code generation studies in subsection II-A. It then provides the relevant LLM for code generation in subsection II-B, leading to the goal of this study.
II-A Code Generation
The task of code generation has evolved significantly over time, and researchers have developed many solutions [15, 16, 17, 18, 19]. Early methods can quickly generate simple code structures based on predefined templates [20] or rule systems [21]. Syriani et al. [17] categorized template-based code generation approaches, identifying their prevalent characteristics and trends, emphasizing the technique’s significance in generating source code across diverse application areas. These simple methods are intuitive to implement and control. However, they have difficulties in implementing complex code and generating code patterns not explicitly defined in the rules [22].
As deep learning (DL) techniques evolved, researchers leveraged approaches such as long short-term memory (LSTM) [23], to comprehend code semantics. By learning from numerous code data, e.g., from the open-source community GitHub, these models can capture the structural and semantic patterns of programming languages, and generate more complex code [24]. Yin et al. [16] introduced a new neural architecture that incorporates a grammar model to explicitly integrate the target syntax as prior knowledge. Their experiments demonstrate that this method is an effective strategy for scaling up to the generation of complex programs from descriptions in natural language. Generally, the DL-based models can produce syntactically correct code, but the generated code was commonly not fully aligned with the given task description [25].
In recent years, the emergence of large language models (LLMs) like GPT [26] and Codex [13] have introduced a new phase in code generation. These models are based on the Transformer architecture [27] and pre-trained on large-scale code datasets to learn the deep semantics and syntactic rules of programming languages. LLMs can better understand developers’ programming requirements in natural language and generate more satisfactory code compared to existing DL-based models [10]. Alawida et al. [28] pointed out that LLMs like ChatGPT strongly influenced the software industries in terms of their architecture, training data, evaluation metrics, etc. Moreover, Zheng et al. [9] reviewed state-of-the-art LLMs and compared their performance for code generation. Unfortunately, they did not uncover why these LLMs failed in code generation.
II-B LLM for Code Generation
With the rise of LLMs exemplified by ChatGPT, many researchers have conducted studies on LLMs-based code generation [29, 30, 31]. These studies encompass various aspects. Previously, many studies focused on building a high-quality evaluation dataset for code generation. To address this issue, Lu et al. [29] introduced CodeXGLUE, a benchmark with 14 datasets to foster DL-based research for program understanding and generation. However, this dataset used evaluation measures like BLEU [32], which cannot verify whether an LLM can generate functionally correct code. Li et al. [33] proposed an open-source large-scale code generation dataset TACO which includes rigorous data processing and unit test validation, although this dataset only focused on competition-level programming questions. Chen et al. [13] proposed HumanEval, an innovative collection of 164 real-world programming challenges in Python. The evaluation encompasses not only syntax but also logical correctness and problem-solving efficiency, requiring the generated code to pass many test cases. HumanEval has made this dataset a widely used benchmark for code generation evaluation [13].
Based on the above evaluation datasets, researchers conducted studies on the factors affecting LLMs’ performance in code generation [34, 35, 36]. Liu et al. [10] found that the performance is sensitive to the LLMs’ prompts, namely the programming requirement written in natural language as model input. By leveraging a chain-of-thought strategy, the optimized prompts can lead to better code generation performance. White et al. [11] focused on improving interactions with LLMs through efficient prompt patterns. They found that prompt patterns significantly enrich the capabilities that can be created in a conversational LLM and the prompt patterns are generalizable to many different domains. Chen et al. [12] introduced Imitation Learning from Language Feedback (ILF). They have shown that ILF can significantly improve the quality of a code generation model, even with just a small sample of human-written feedback and refinements. Diao et al. [37] proposed a new method, Active-Prompt, to adapt LLMs to different tasks with task-specific example prompts. Active-Prompt achieved a promising performance on eight widely used datasets for arithmetic reasoning, commonsense reasoning, and symbolic reasoning. Zhao et al. [38] proposed Logical Thoughts (LoT) to improve the zero-shot chain-of-thought reasoning ability of large language models. Their experiments demonstrate that the method leads to improved reasoning ability, particularly when applied to LLMs of large scale. In summary, existing studies mainly focus on prompt optimization and have not investigated why LLMs did not pass the test cases. To fill this gap, this study aims to analyze why LLMs failed and explore the way to directly fix the generated code.
III Empirical Study
III-A Research Questions
This empirical study aims to systematically analyze and compare the performance of various LLMs, and to categorize the errors in the generated results to guide potential fixing methods. To achieve this goal, we investigate the following RQs:
RQ1: Can we reproduce the performance of state-of-the-art LLMs? This RQ examines the reproducibility of results claimed in the literature on LLMs for code generation. By attempting to replicate the conditions and methodologies of previous studies, we aim to assess the validity of reported performance metrics. This RQ is crucial for establishing a baseline of expectations when it comes to the operational capabilities of LLMs.
RQ2: Why did these LLMs fail in code generation? Investigating the code generation errors of LLMs is essential for limitation understanding and further improvements. By running the generated code with test cases and analyzing the occurred errors, we aim to uncover the underlying factors that contribute to these errors.
III-B Dataset and Evaluation Measure
To conduct the empirical study, we utilize the widely used dataset HumanEval [13], meticulously crafted by OpenAI, for assessing various LLMs. The HumanEval dataset is an innovative collection of 164 real-world programming challenges in Python. Each generation task involves an average of 7.7 test cases, ensuring the functional correctness of the generated code. Most of the well-known LLMs such as CodeLlama [19], CodeGen [39], and WizardCoder [40] utilized HumanEval for evaluating code generation capabilities. As shown in Fig. 1, the HumanEval comprises five components: 1) task_id, a unique identifier for each task, allowing for easy reference and organization of the code generation tasks. 2) prompt, a textual statement that provides generation requirements for guiding LLMs. 3) canonical_solution, the standard solution for each task. 4) test, a testing function along with several test cases, designed to meticulously assess whether the generated code meets the expected requirement. 5) entry_point, the name of a function, which is intended to be the main entry point of the generated code function.
Figure 1: A schematic diagram of the first sample in HumanEval.
For this dataset, an LLM needs to generate the functionally correct code, which can pass all the test cases, according to a given prompt. To measure the accuracy of code generation, Chen et al. [13] provide the $Pass@k$ metric, the percentage of tasks solved by an LLM. A task is considered solved if any top-$k$ code generated by LLM can pass all the test cases. As Pass@k has high variance, we follow Chen et al. [13] and use the unbiased version: $Pass@k=\mathbb{E}\left[1-\binom{n-c}{k}/\binom{n}{k}\right]$, where $\mathbb{E}$ denotes the average performance across all tasks; $n$ is the total number of tasks; $c$ is the number of correctly solved tasks; $\binom{n}{k}$ is the number of combinations for choosing $k$ tasks out of $n$; $\binom{n-c}{k}$ is the number of combinations for choosing $k$ tasks from those one model cannot solve (i.e., $n-c$ tasks), representing the number of ways to fail. Note that this study measured the code generation performance with $Pass@1$, as developers commonly expected that LLMs could generate correct code in the first place. Besides, to mitigate the effects of randomness in LLMs, we follow [41] to test each LLM ten times and then report the average value as the final $Pass@1$ result.
III-C Can We Reproduce the Performance of State-of-the-Art LLMs? (RQ1)
III-C1 Motivation and Method
In recent years, researchers have developed many LLMs and published their test results on the HumanEval dataset [9]. However, achieving outcomes that are completely identical to those reported remains a challenging task [42]. Therefore, to thoroughly investigate whether existing LLMs are consistent with the reported performances as described in RQ1, we adopted several state-of-the-art LLMs for testing. One is the ChatGPT, a state-of-the-art closed-source LLM released by OpenAI [43]. Under zero-shot conditions, ChatGPT has achieved an accuracy rate of over 60% on the HumanEval dataset maintaining a leading position [44].
For open-source LLMs, Zheng et al. [9] reported on the performance of many LLMs on HumanEval. Based on this study, we selected four families of LLMs. Phind [45] performs best in the reported ranking, followed by WizardCoder [40] and CodeLlama [19]. Phi [46] is the best small-sized LLM and is competitive with other families. Besides, we included the CodeTulu family [47] in our study due to its promising performance as reported by Ivison et al. [47]. All open-source LLMs are re-implemented by using the LLM parameters stored in the HuggingFace platform and tested on a server with four NVIDIA RTX3090 GPUs in default settings.
Detailed information about each LLM family and relevant links are listed below. We ran each LLM ten times on HumanEval and compared the code generation performance from original reports [9, 47].
@@ III-C1 Motivation and Method
    ChatGPT, is a powerful LLM released by OpenAI [48]. Researchers have demonstrated its high performance in code generation [49]. We chose the GPT-3.5-turbo version with 175B parameters by invoking official APIs [50] following Sun et al. [51].

### v2

# Fixing Function-Level Code Generation Errors for Foundation Large Language Models

@@
Abstract.
Function-level code generation leverages foundation Large Language Models (LLMs) like CodeLlama to automatically produce source code with expected functionality. It has been widely investigated and applied in intelligent programming assistants, such as GitHub Copilot, to enhance software development productivity and reduce repetitive tasks. Despite advancements in foundation LLMs, the generation involves many errors. Existing studies leverage static analysis tools (e.g., TBar and ReCoder) or add another fixing LLM (i.e., CoR and LDB) to post-process these errors. However, there are still many errors remaining to be solved because their root causes have not been investigated yet, making it challenging to design better fixing tools.
@@ Abstract.
In this paper, we first conducted an empirical study on the generation errors. Specifically, we reproduced 14 representative open/closed source LLMs on the HumanEval dataset and verified their correctness. We obtained 12,837 code generation errors and conducted an in-depth analysis of their causes, leading to 19 categories of error causes. Our empirical analysis indicated that three of these causes can be directly fixed. Based on the findings, we proposed a fixing method called LlmFix, which is a rule-based static analysis approach and addresses these three types of errors through a three-step process: filtering code for indentation correction, truncating redundant generated code, and importing missing modules. Evaluations of LlmFix are conducted from two perspectives: its performance on error-fixing tasks and its impact on improving function-level code generation tasks. For error fixing performance, we built an evaluation dataset LlmErrorEval based on our collected 12k+ errors. Experimental results show that LlmFix achieves a fix rate of 17.1% outperforming the best state-of-the-art LDB by 8.9%. For code generation improvements, evaluations of LlmFix on both the HumanEval and MBPP datasets demonstrate its effectiveness, significantly improving code generation accuracy by an average Pass@1 of 7.5% across 14 LLMs on HumanEval and MBPP.
@@
Foundation Large Language Model, Function-Level Code Generation, Automatic Error Fixing
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 8ccs: Software and its engineering Software development techniques
1. Introduction
Code generation refers to automatically converting natural language descriptions of software development specifications into code snippets written in a programming language with correct functionality (Manning, 2003; Gabriel et al., 2022). Intelligent programming assistants like GitHub Copilot (Chen et al., 2021) can help developers reduce coding efforts in writing repetitive code functions and improve their development productivity (Frank et al., 2020; Guo et al., 2020). Essentially, these assistants are powered by Large Language Models (LLMs), such as Codex (Chen et al., 2021) and CodeLlama (Roziere et al., 2023). These foundation models are pre-trained on extensive codebases and fine-tuned specifically for code generation tasks. To evaluate their performance, researchers have developed various benchmarks, which can be categorized into function-level (e.g., HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)), class-level (e.g., ClassEval (Du et al., 2023)), and repo-level (e.g., Swe-bench (Jimenez et al., 2024)). Function-level benchmarks focus on the correctness of code functionality, while class-level and repo-level benchmarks evaluate code generation that requires the generated code depending on existing APIs implemented in the project for necessary reuse consideration. In this study, we focus on the correctness of code functionality, namely function-level code generation, which is also widely investigated by researchers (Chen et al., 2021; Austin et al., 2021) and applied in intelligent programming assistants (Wong et al., 2023; Corso et al., 2024).
Despite advancements in foundation LLMs, there are still many errors even in function-level code generation (Hong et al., 2024; Huang et al., 2024). To address these errors, Fan et al. (Fan et al., 2023) investigated the viability of fixing them by leveraging existing static analysis tools, e.g., TBar (Liu et al., 2019) and ReCoder (Zhu et al., 2021), but their experimental results indicated that the fixing capability of these tools is ineffective. Recently, researchers tried to add another fixing LLM to address the errors produced by the foundation LLMs, such as Large Language Model Debugger (LDB) (Zhong et al., 2024) and Chain of Repair (CoR) (Wang et al., 2023). Although these methods show better effectiveness, there are still many errors remaining to be solved (Atil et al., 2024). This is because existing studies have not systematically investigated the root causes of function-level code generation errors in foundation LLMs, making it difficult to develop more effective error-fixing tools.
In this paper, we first conduct an empirical study to evaluate the performance of existing foundation LLMs and identify errors that can be directly fixed. Specifically, we reproduced 14 representative open/closed-source LLMs (e.g., CodeLlama-Python) and verified the reproduction correctness. The experimental results demonstrate that these 14 LLMs exhibit strong reproducibility in performance. Subsequently, we analyzed the root cause of 12,837 errors produced by these 14 LLMs and classified them into 19 categories, suggesting a variety of code generation errors. Besides, we investigated the repairability of these errors according to the given limited context. We found that three categories can be directly fixed, including missing import, redundant generation, and inconsistent indentation.
@@ 1. Introduction
Based on the empirical findings, we design LlmFix, a rule-based static analysis method specifically tailored to automatically fix the three repairable error categories. Generally, LlmFix works in three steps: 1) Code Filtering, searching for inconsistent indentation, and replacing it with uniform tab-based indentation. 2) Code Truncation, discarding redundant code located after the function completion. 3) Importing Missing Modules, automatically adding corresponding import statements at the beginning of the code when detecting the use of a module that has not been imported.
We conducted a comprehensive evaluation of LlmFix from the following two perspectives: 1) Error Fixing Performance. We constructed a dataset, LlmErrorEval, for error-fixing tasks based on the collected 12k+ errors to evaluate the performance of existing fixing methods. Experimental results show that LlmFix achieves a fix rate of 17.1%, outperforming the best baseline LDB by 8.9%. The fix rate can be further enhanced to 24.9% by integrating LlmFix with LDB. The integration brings little extra computation overhead as LlmFix only takes 11.5ms to run. 2) Code Generation Improvements. We applied LlmFix to 14 LLMs for code generation tasks. Experimental results showed that LlmFix is effective for all LLMs and the generation accuracy can be significantly improved by 7.5% on average.
@@
  • Conducting a systematic investigation into the performance of foundation LLMs on function-level code generation tasks, categorizing 14 types of errors and analyzing their distributions and 19 types of root causes.
  • Proposing a rule-based static analysis method, LlmFix (Hao et al., 2024), to automatically fix three identified fixable errors: missing imports, redundant generation, and inconsistent indentation.
  • Constructing a dataset, LlmErrorEval (Hao et al., 2024), comprising 12,837 data samples to evaluate the effectiveness of error-fixing methods.
This paper is structured as follows. Section 2 introduces the related work on code generation. Section 3 provides the empirical study on LLM reproduction and error analysis. Section 4 presents our error fixing method LlmFix, which is evaluated in Section 5 and discussed in Section 6. Section 7 summarizes this paper.
2. Related Work
2.1. Function-Level Code Generation
The task of code generation has evolved significantly over time, and researchers have developed many solutions (Puschel et al., 2005; Pengcheng and Graham, 2017; Syriani et al., 2018; Svyatkovskiy et al., 2020; Roziere et al., 2023). Early methods can quickly generate simple code structures based on predefined templates (Danilchenko and Fox, 2012) or rule systems (Bajwa et al., 2006). Syriani et al. (Syriani et al., 2018) categorized template-based code generation approaches, identifying their prevalent characteristics and trends, emphasizing the technique’s significance in generating source code across diverse application areas. These simple methods are intuitive to implement and control. However, they have difficulties in implementing complex code and generating code patterns not explicitly defined in the rules (van der Lee et al., 2018).
As deep learning (DL) techniques evolved, researchers leveraged approaches such as long short-term memory (LSTM) (Memory, 2010), to comprehend code semantics. By learning from numerous code data, e.g., from the open-source community GitHub, these models can capture the structural and semantic patterns of programming languages, and generate more complex code (Le et al., 2022). Yin et al. (Pengcheng and Graham, 2017) introduced a new neural architecture that incorporates a grammar model to explicitly integrate the target syntax as prior knowledge. Their experiments demonstrate that this method is an effective strategy for scaling up to the generation of complex programs from descriptions in natural language. Generally, the DL-based models can produce syntactically correct code, but the generated code was commonly not fully aligned with the given task description (Yang et al., 2023).
In recent years, the emergence of LLMs like ChatGPT (Brown et al., 2020) and CodeLlama (Roziere et al., 2023) have introduced a new phase in code generation. These models are based on the Transformer architecture (Vaswani et al., 2017) and pre-trained on large-scale code datasets to learn the deep semantics and syntactic rules of programming languages. ChatGPT is a representative of closed-source LLMs, which have been widely used and possess strong code generation capabilities (60.30% on HumanEval) (OpenAI, 2024a). In the realm of open-source large models, Rozière et al. (Roziere et al., 2023) introduced the CodeLlama series, based on the LLAMA2 architecture, designed to support long context inputs specifically for code generation tasks. Ivison et al. (Ivison et al., 2023) proposed the CodeTulu-2 series, which is based on the CodeLlama model and fine-tuned on a series of high-quality instruction datasets established by them, named TÜLU-V2-mix, achieving better performance than the CodeLlama model. Luo et al. (Luo et al., 2023) applied the Evol-Instruct method to the code domain, introducing the WizardCoder series, which enables complex instruction fine-tuning. Gunasekar et al. (Gunasekar et al., 2023) utilized “textbook quality” data from the internet and synthesized textbooks and exercises using ChatGPT, training the Phi series models characterized by a small parameter count (1.3B-2.7B) and high code generation performance (50.6% on HumanEval).
2.2. Error Fixing for Foundation LLMs
Austin et al. (Austin et al., 2021) indicated that existing foundation LLMs involve many generation errors remaining to be solved. There are two types of methods for fixing generation errors.
The first type is the static analysis method. These methods typically rely on program semantics or syntax for code error fixing. Specifically, Smytzek et al. (Smytzek et al., 2024) proposed a recent code error fixing framework called FixKit, which combines five widely used code error fixing methods: 1) GENPROG (Weimer et al., 2009): A genetic programming-based approach that evolves program variants to fix bugs by applying mutation operators to the abstract syntax tree of the program; 2) MUTREPAIR (Debroy and Wong, 2010): A mutation-based fixing technique that generates patches by applying mutation operators to the faulty program; 3) KALI (Qi et al., 2015): A fixing technique that focuses on removing or modifying potentially faulty code to generate patches; 4) CARDUMEN (Martinez and Monperrus, 2018): An automatic program fixing technique that uses a probabilistic model to generate fix templates based on mined code patterns; 5) AE (Weimer et al., 2013): An adaptive search-based technique that uses program equivalence relations to guide the patch generation process. This framework makes these methods easier to use and compare in terms of performance. However, the authors did not evaluate the effectiveness of these methods in fixing errors in LLM-generated code, which we have addressed in section 5. Fan et al. (Fan et al., 2023) summarized the characteristics of automatically generated code and evaluated two fixing methods: TBar (Liu et al., 2019) and Recoder (Zhu et al., 2021). TBar represents a search-based and pattern-based fixing tool, whereas Recoder is a learning-based approach. Their experiments showed that these two methods are ineffective in fixing errors in generated code: they have limited search spaces, are unable to generate multi-edit patches, and lack awareness of program dependencies. However, the authors did not propose new effective static analysis methods, which is precisely the focus of our research.
The second type is the fixing LLM. The basic idea of this type is to build another fixing LLM to post-process the code generation errors produced by the foundation LLMs. Specifically, Zhong et al. (Zhong et al., 2024) introduced the Large Language Model Debugger (LDB) method, an LLM-based fixing approach that allows LLMs to use runtime execution information to improve generated programs and fix errors. Wang et al. (Wang et al., 2023) introduced CoR, a method designed to emulate the human process of debugging. It works by iteratively gathering feedback from compilers, interpreting the causes of errors and crafting solutions in natural language, and ultimately applying fixes to the program.
3. Empirical Study
3.1. Research Questions
This empirical study aims to evaluate the reproducibility of existing representative foundation LLMs, identify the root causes of code generation errors produced by these LLMs, and determine which errors are repairable to guide potential fixing methods.
To achieve this goal, we investigate the following RQs:
RQ1: Can we reproduce the performance of existing foundation LLMs? This RQ examines the reproducibility of results claimed in the literature on LLMs for code generation. By attempting to replicate the conditions and methodologies of previous studies, we aim to assess the validity of reported performance metrics. This RQ is crucial for establishing a baseline of expectations when it comes to the operational capabilities of LLMs.
RQ2: Why did these foundation LLMs fail in code generation? Investigating the code generation errors of LLMs is essential for limitation understanding and further improvements. By running the generated code with test cases and analyzing the occurred errors, we aim to uncover the underlying factors that contribute to these errors.
3.2. Dataset and Evaluation Measure
To conduct the empirical study, we utilize the widely used dataset HumanEval (Chen et al., 2021), meticulously crafted by OpenAI, for assessing various LLMs. The HumanEval dataset is a collection of 164 real-world programming challenges. Each generation task involves an average of 7.7 test cases, ensuring the functional correctness of the generated code. Most of the well-known LLMs such as CodeLlama (Roziere et al., 2023), CodeTulu (Ivison et al., 2023), and WizardCoder (Luo et al., 2023) utilized HumanEval for evaluating code generation capabilities. The HumanEval comprises five components: 1) task_id, a unique identifier for each task, allowing for easy reference and organization of the code generation tasks. 2) prompt, a textual statement that provides generation requirements for guiding LLMs. 3) canonical_solution, the standard solution for each task. 4) test, a testing function along with several test cases, designed to meticulously assess whether the generated code meets the expected requirement. 5) entry_point, the name of a function, which is intended to be the main entry point of the generated code function.
For this dataset, an LLM needs to generate the functionally correct code, which can pass all the test cases, according to a given prompt. To measure the accuracy of code generation, Chen et al. (Chen et al., 2021) provide the $Pass@k$ metric, the percentage of tasks solved by an LLM. A task is considered solved if any top-$k$ code generated by LLM can pass all the test cases. As Pass@k has high variance, we follow Chen et al. (Chen et al., 2021) and use the unbiased version: $Pass@k=\mathbb{E}\left[1-\binom{n-c}{k}/\binom{n}{k}\right]$, where $\mathbb{E}$ denotes the average performance across all tasks; $n$ is the total number of tasks; $c$ is the number of correctly solved tasks; $\binom{n}{k}$ is the number of combinations for choosing $k$ tasks out of $n$; $\binom{n-c}{k}$ is the number of combinations for choosing $k$ tasks from those one model cannot solve (i.e., $n-c$ tasks), representing the number of ways to fail. Note that this study measured the code generation performance with $Pass@1$, as developers commonly expected that LLMs could generate correct code in the first place. Besides, to mitigate the effects of randomness in LLMs, we follow (Liu et al., 2021a) to test each LLM ten times and then report the average value as the final $Pass@1$ result.
3.3. Can We Reproduce the Performance of Existing Foundation LLMs? (RQ1)
3.3.1. Motivation and Method
In recent years, researchers have developed many foundation LLMs and published their test results on the HumanEval dataset (Zheng et al., 2023). However, achieving outcomes that are completely identical to those reported remains a challenging task (Chang et al., 2024). Liu et al. (Liu et al., 2021b) conducted a literature review on DL studies published in twenty prominent SE venues over the past five years. Their findings indicate that DL models inherently possess randomness, which can compromise the replicability of studies involving unstable models. A DL model with a low level of convergence can exhibit highly variable and unstable performance, resulting in poor reproducibility of the studies. Therefore, to thoroughly investigate whether existing LLMs are consistent with the reported performances as described in RQ1, we adopted several LLMs for testing. One is the ChatGPT, a closed-source LLM released by OpenAI (OpenAI, 2024b).
For open-source LLMs, Zheng et al. (Zheng et al., 2023) reported on the performance of many LLMs on HumanEval. Based on this study, we selected four families of LLMs. Phind (Phind, 2023) performs best in the reported ranking, followed by WizardCoder (Luo et al., 2023) and CodeLlama (Roziere et al., 2023). Phi (Gunasekar et al., 2023) is the best small-sized LLM and is competitive with other families. Besides, we included the CodeTulu family (Ivison et al., 2023) in our study due to its promising performance as reported by Ivison et al. (Ivison et al., 2023). All open-source LLMs are re-implemented by using the LLM parameters stored in the HuggingFace platform and tested on a server with four NVIDIA RTX3090 GPUs in default settings.
Detailed information about each LLM family and relevant links are listed below. We ran each LLM ten times on HumanEval and compared the code generation performance from original reports (Zheng et al., 2023; Ivison et al., 2023).
@@ 3.3.1. Motivation and Method
    ChatGPT, is a powerful LLM released by OpenAI (Achiam et al., 2023). Researchers have demonstrated its high performance in code generation (Wu et al., 2023). We chose the GPT-3.5-turbo version with 175B parameters by invoking official APIs (OpenAI, 2024c) following Sun et al. (Sun et al., 2023).
@@
    Phind, was introduced by the Phind team, available in two versions (PhindTeam, 2023): Phind-CodeLlama-34B-v1 and Phind-CodeLlama-34B-v2. According to the team’s report (Phind, 2023), the latter version even outperforms OpenAI’s GPT series in code generation.
@@
    WizardCoder, was introduced by Microsoft (Luo et al., 2023), which empowers code generation LLMs with complex instruction fine-tuning. We re-implemented the versions of WizardCoder-Python in different parameter sizes (7B/13B/34B) (Microsoft, 2023b).
@@
    CodeLlama, was introduced by Meta AI (Roziere et al., 2023), which is a competitive code generation model and inspired many other varied versions such as CodeTulu (Ivison et al., 2023) and WizardCoder (Luo et al., 2023). We used the CodeLlama-python (7B/13B/34B) (MetaAI, 2023) as our baseline, which fine-tunes CodeLlama with more Python code and shows better performance in code generation.
@@
    Phi, was released by Microsoft (Gunasekar et al., 2023), which is renowned for its smaller model size (1.3B) but demonstrated exceptional high performance in code generation. Later, Microsoft released new versions including Phi-1.5 Li et al. (Li et al., 2023) and Phi-2 (Microsoft, 2023c). Thus, these three models were selected as our baselines including Phi-1 (1.3B), Phi-1.5 (1.3B), and Phi-2 (2.7B) (Microsoft, 2023a).